# A Comprehensive Survey on Knowledge Distillation

### v1

@@ VI-B2 Segment Anything (SAM)
SAM is utilized for weakly supervised semantic segmentation [571, 572, 573, 574, 575] by leveraging its pre-trained capabilities to generate high-quality pseudo-labels from minimal or weak supervision signals, such as image-level tags, points, or bounding boxes. These pseudo-labels serve as a foundation for training more efficient models, bridging the gap between limited supervision and fully labeled data. The robustness of SAM’s outputs allows semantic segmentation models to learn complex patterns and object boundaries without extensive manual annotation. This approach significantly reduces labeling costs while maintaining competitive performance.
VI-B3 Multi-model Applications
In some instances, research has explored synergistic use of multiple FMs, combining their strengths through KD to enhance performance in segmentation, detection, and multi-modal applications. This collaborative use of KD fosters the development of unified models that inherit the complementary strengths of various teacher architectures. General studies have demonstrated the effectiveness of integrating diverse FMs across tasks [576, 577, 578, 579, 580, 581, 582, 583], highlighting improvements in both efficiency and accuracy. One common approach involves combining models with complementary strengths. For instance, (SAM + CLIP) [584, 585, 586] leverages SAM’s robust image segmentation alongside CLIP’s semantic understanding, enhancing performance in various vision tasks. Similarly, (DINO + CLIP) [587, 588] has been explored, taking advantage of DINO’s self-supervised learning capabilities and CLIP’s strong vision-language alignment. Extending this idea, (DETR + CLIP) has been proposed in [589] for open-vocabulary object detection, where DETR serves as the detection model while CLIP provides text-based prompts. Additionally, research [590, 591] has investigated using the complementary strengths of (SAM + GDINO) for object detection and segmentation.
VI-B4 Vision Transformers
@@
ViT [494] is a deep learning model that extends the Transformer architecture [592], originally developed for natural language processing, to image analysis. Several works [593, 594, 595, 596, 597, 598, 599] have explored the application of KD to ViT, investigating a range of strategies aimed at achieving diverse objectives, as summarized in Table IX.
TABLE IX: Summary of Methods Applying Knowledge Distillation to the Vision Transformer Model.
@@
SMKD [593] (2023) Explores the tendency of ViTs to overfit and degrade in performance under few-shot learning, due to the lack of CNN-like inductive biases, and proposes a KD method to address this issue.
TinyMIM [594] (2023) Explores KD techniques for transferring the benefits of large MIM (Masked Image Modeling) pre-trained models to smaller ones.
Incrementer [595] (2023) Introduces a class-incremental semantic segmentation framework that employs ViT instead of traditional CNNs, emphasizing the need for global context modeling and incorporating convolutional layers to support new class predictions in incremental learning settings.
CST [596] (2023) Uses self-distillation by leveraging attention maps from a frozen ViT to generate pseudo-labels for few-shot, weakly supervised tasks.
DMAE [597] (2023) Suggests distilling intermediate features, rather than logits, from a large ViT model to a smaller one.
CSKD [598] (2023) Presents a method for enhancing ViT performance through KD from CNN models.
ViTKD [599] (2024) Introduces a method for distilling knowledge between ViT feature maps instead of logits, using DeiT [497] as the base model.
@@
DINO [493] is a self-supervised vision transformer that learns rich visual representations from unlabeled data. Due to its strong ability to capture meaningful features, many works leverage Dino’s pretrained features for object localization for various downstream tasks, such as unsupervised object discovery [600, 601, 602, 603, 478, 604], open-world counting [605], and zero-shot semantic segmentation [606].
@@
DETR [495] is an end-to-end object detection model built on ViT, capable of directly predicting bounding boxes and class labels within a unified framework, thereby eliminating the need for anchors or complex post-processing. Its ability to effectively capture global context and model interactions between objects has positioned it as a leading backbone in object detection tasks. However, DETR’s large model size and substantial computational demands pose significant challenges for deployment in real-world scenarios with limited computational budgets. To overcome these limitations, recent works [607, 608] have explored compressing DETR through KD, aiming to retain its strong performance while reducing computational overhead for real-time and resource-limited applications.
@@
Swin Transformer [496] is a powerful architecture that effectively combines the strengths of CNNs and ViTs through its hierarchical design and efficient window-based attention mechanism. By employing local window attention within each window and shifting the windows between layers, Swin captures both local and global context, enabling it to excel in modeling fine-grained details as well as long-range dependencies. This rich feature representation has made Swin a preferred choice in various computer vision tasks. Recent works [609, 610, 611, 612, 613, 614, 615, 616, 617, 618, 619, 620, 621] have leveraged Swin Transformer as an encoder, utilizing its features across different vision tasks, as summarized in Table X.
TABLE X: Summary of Methods Applying Knowledge Distillation to the Swin Transformer Model.
@@
MaskFormer [611] (2021) Semantic Segmentation
DFR [610] (2021) Weakly Supervised Semantic Segmentation (Scribbles)
SwinNet [612] (2021) Salient Object Detection
SwinTrack [613] (2022) Siamese Tracking
SwinF [615] (2022) Target Detection
Swin-UNETR [616] (2022) Medical Image Analysis
Ssformer [614] (2022) Semantic Segmentation
HGI-SAM [617] (2023) Weakly Supervised Intracranial Hemorrhage Segmentation
HiFormer [618] (2023) Medical Image Segmentation
Swin-Fusion [619] (2023) Human Action Recognition
P.Swin [620] (2023) Image Classification and Object Detection
SCUNet++ [621] (2024) Pulmonary Embolism CT Image Segmentation
@@
SWTformer [609] (2024) Weakly Supervised Semantic Segmentation (Class Labels)
VI-C Distillation in Self-supervised learning
@@
Self-Supervised Learning (SSL) is an important field due to its independence from labeled data, and self-distillation is the basis of the most prominent works in SSL, especially in the realm of computer vision. The most prominent work in SSL was based on self-distillation [622]. SimCLR [622] proposed a shared network that receives two different augmentations of an image and minimizes their extracted embeddings. Following SimCLR, different extensions have been proposed [623, 624, 625, 626, 627, 628]. MOCO [623, 624] and BYOL [625] are among successful methods where they propose to use the same network but with a different training process. The student is trained using backpropagation, and the teacher is an EMA of the student. SwAP [627] uses some prototypes and assigns a code to the augmentations of the same image. The student is then trained to predict the same class ID as the teacher. DINO [493] and DINOv2 [629] are the most recent important applications of KD in SSL, where global and local patches of an image pass through similar teacher and student networks. The teacher is updated using EMA, and the representations of the teacher and the student are forced to be similar. This local-to-global approach makes it ideal for segmentation tasks.
@@
Recently, more complex forms of SSL based on KD have been proposed [626, 628, 444, 630, 631, 632]. BINGO [628] uses a pre-trained SSL method to group the images of a dataset into a bag of samples. Then, in each epoch, samples are taken from the bag and the inter and intra-sample losses are minimized. [444] trains two SSL teachers for images and videos using masked modeling and follows a scenario similar to SimCLR, minimizing the spatial features with the image teacher and the spatio-temporal features with the video teacher. [630] uses a parallel SSL approach and defines consistency losses between them, and [631] integrates a known KD approach into an SSL framework using an SSL pre-trained teacher.
@@
Apart from the mentioned papers, some studies attempt to distill the knowledge of a pre-trained SSL network into a smaller network [633, 634, 635, 636, 637]. The gap between two similar networks, one trained fully-supervised and one trained self-supervised, is shown to decrease with a bigger model size [633]. [635] shows the impact of an MLP layer after representations of the student, while [633] proposes a similarity-based distillation method for distilling the knowledge of a pre-trained network, which can outperform the corresponding fully-supervised trained network. In a similar approach, SEED [636] uses a queue of samples and minimizes the distance of the student’s representation with anchors in the queue and the corresponding distances between the anchors and the teacher’s representation. [637] proposes a novel approach to complement self-supervised pretraining via an auxiliary pretraining phase for small unlabeled datasets.
VI-D Diffusion Distillation
@@
Diffusion models [73, 638] generate high-quality images, text, and 3D data through an iterative denoising process that can require hundreds or thousands of steps. This is computationally expensive, limiting deployment on low-resource hardware. KD mitigates this by transferring soft outputs and internal representations from a large teacher model to a smaller student model. The distilled model then achieves the same quality with far fewer sampling steps, requiring significantly less inference time and computation [639].
@@
Several strategies have been developed to accelerate diffusion models. Feature-level methods have the student mimic the teacher’s internal representations [72, 640, 641, 642]. Sampling process distillation reduces the iterative denoising by either merging multiple steps into one or matching a one-step student’s output to that of the full teacher [643, 644]. Adversarial and data-free approaches align outputs using GAN losses or bootstrapping without requiring large synthetic datasets [160, 645, 646]. Consistency models reformulate the denoising into a direct, often single-step transformation, cutting down inference time [647, 648, 649, 650, 651, 652]. In the following sections, each category is explained in detail.
@@
Feature-level: Feature-level distillation accelerates diffusion model efficiency by bringing the student representations closer to the teacher and reducing inference steps. Recent methods reinforce feature alignment through denoising [72], external guidance [640], and cross-sample consistency [641], which facilitate convergence and improved sampling efficiency. Furthermore, combining feature distillation with model compression reduces model size and inference time [642]. These techniques accelerate generation without compromising output quality, thereby making diffusion models more realistic.
@@
Sampling Process: Distillation of the sampling procedure is performed to ease the naturally iterative denoising process characteristic of diffusion models. Progressive distillation [643] achieves this by consolidating two steps into a single step; through repeated application of this simplification, the duration of the sampling procedure can be drastically shortened. Distribution matching distillation [644] attempts to condense the entire iterative procedure into a single step. This alignment is accomplished by matching the output distribution of a one-step student model directly with that of the combined multi-step teacher model.
@@
Adversarial and Data-free: Adversarial methods [160, 645] leverage GAN-based losses to enforce the student’s output distribution to closely match that of the teacher. Alternatively, data-free methods like EM distillation [646], employ bootstrapping between successive denoising steps, eliminating the need for large synthetic datasets during training.
@@
Consistency Models: Consistency models [647, 648, 649] reformulate the denoising process as a direct, often single transformation from noisy data to its clean version. Their latent-space variants [650, 651], along with improvements focusing on stability and scalability [652], achieve significant speedups in inference times. However, training these models may require careful tuning to prevent quality collapse.
@@
Furthermore, recent developments in the area of distillation techniques expanded their accessibility to multi-modal and multi-model settings, thus improving transferability across tasks and diverse architectures. ProlificDreamer [653] and DREAMFUSION [654] utilize the methods in the realm of text-to-3D translation, and general-purpose systems like Diff-Instruct [655] can facilitate generalization for the task.
@@
VI-E Visual Recognition Distillation

### v2

@@ 6.2.2 Segment Anything (SAM)
SAM is utilized for weakly supervised semantic segmentation (Kweon & Yoon, 2024; Jiang & Yang, 2023; Chen & Sun, 2023; Chen et al., 2023c; Sun et al., 2023a) by leveraging its pre-trained capabilities to generate high-quality pseudo-labels from minimal or weak supervision signals, such as image-level tags, points, or bounding boxes. These pseudo-labels serve as a foundation for training more efficient models, bridging the gap between limited supervision and fully labeled data. The robustness of SAM’s outputs allows semantic segmentation models to learn complex patterns and object boundaries without extensive manual annotation. This approach significantly reduces labeling costs while maintaining competitive performance.
6.2.3 Multi-model Applications
In some instances, research has explored synergistic use of multiple FMs, combining their strengths through KD to enhance performance in segmentation, detection, and multi-modal applications. This collaborative use of KD fosters the development of unified models that inherit the complementary strengths of various teacher architectures. General studies have demonstrated the effectiveness of integrating diverse FMs across tasks (Yin & Jiang, 2024; Ranzinger et al., 2024; Sun et al., 2023c; Ye et al., 2023a; Shang et al., 2024; Gao et al., 2024; Englert et al., 2024; Rai et al., 2024), highlighting improvements in both efficiency and accuracy. One common approach involves combining models with complementary strengths. For instance, (SAM + CLIP) (Wang et al., 2024d; Aleem et al., 2024; Yang & Gong, 2024) leverages SAM’s robust image segmentation alongside CLIP’s semantic understanding, enhancing performance in various vision tasks. Similarly, (DINO + CLIP) (Wysoczańska et al., 2024b; Kerr et al., 2023) has been explored, taking advantage of DINO’s self-supervised learning capabilities and CLIP’s strong vision-language alignment. Extending this idea, (DETR + CLIP) has been proposed in Jiang et al. (2024b) for open-vocabulary object detection, where DETR serves as the detection model while CLIP provides text-based prompts. Additionally, research (Ren et al., 2024; Das et al., 2024) has investigated using the complementary strengths of (SAM + GDINO) for object detection and segmentation.
6.2.4 Vision Transformers
@@
ViT (Dosovitskiy, 2020) is a deep learning model that extends the transformer architecture (Vaswani, 2017), originally developed for natural language processing, to image analysis. Several works (Lin et al., 2023a; Ren et al., 2023a; Shang et al., 2023a; Kang et al., 2023b; Bai et al., 2023; Zhao et al., 2023a; Yang et al., 2024) have explored the application of KD to ViT, investigating a range of strategies aimed at achieving diverse objectives, as summarized in Table 9.
Table 9: Summary of Methods Applying Knowledge Distillation to the Vision Transformer Model.
@@
SMKD (Lin et al., 2023a) Explores the tendency of ViTs to overfit and degrade in performance under few-shot learning, due to the lack of CNN-like inductive biases, and proposes a KD method to address this issue.
TinyMIM (Ren et al., 2023a) Explores KD techniques for transferring the benefits of large MIM (Masked Image Modeling) pre-trained models to smaller ones.
Incrementer (Shang et al., 2023a) Introduces a class-incremental semantic segmentation framework that employs ViT instead of traditional CNNs, emphasizing the need for global context modeling and incorporating convolutional layers to support new class predictions in incremental learning settings.
CST (Kang et al., 2023b) Uses self-distillation by leveraging attention maps from a frozen ViT to generate pseudo-labels for few-shot, weakly supervised tasks.
DMAE (Bai et al., 2023) Suggests distilling intermediate features, rather than logits, from a large ViT model to a smaller one.
CSKD (Zhao et al., 2023a) Presents a method for enhancing ViT performance through KD from CNN models.
ViTKD (Yang et al., 2024) Introduces a method for distilling knowledge between ViT feature maps instead of logits, using DeiT (Touvron et al., 2021) as the base model.
@@
DINO (Caron et al., 2021) is a self-supervised vision transformer that learns rich visual representations from unlabeled data. Due to its strong ability to capture meaningful features, many works leverage Dino’s pretrained features for object localization for various downstream tasks, such as unsupervised object discovery (Kara et al., 2024; Siméoni et al., 2021; 2023; Wang et al., 2023n; 2022b; Liu et al., 2024f), open-world counting (Amini-Naieni et al., 2024), and zero-shot semantic segmentation (Karazija et al., 2023).
@@
DETR (Carion et al., 2020) is an end-to-end object detection model built on ViT, capable of directly predicting bounding boxes and class labels within a unified framework, thereby eliminating the need for anchors or complex post-processing. Its ability to effectively capture global context and model interactions between objects has positioned it as a leading backbone in object detection tasks. However, DETR’s large model size and substantial computational demands pose significant challenges for deployment in real-world scenarios with limited computational budgets. To overcome these limitations, recent works (Chang et al., 2023; Wang et al., 2024m) have explored compressing DETR through KD, aiming to retain its strong performance while reducing computational overhead for real-time and resource-limited applications.
@@
Swin Transformer (Liu et al., 2021e) is a powerful architecture that effectively combines the strengths of CNNs and ViTs through its hierarchical design and efficient window-based attention mechanism. By employing local window attention within each window and shifting the windows between layers, Swin captures both local and global context, enabling it to excel in modeling fine-grained details as well as long-range dependencies. This rich feature representation has made Swin a preferred choice in various computer vision tasks. Recent works (Ahmadi & Kasaei, 2024; Zhang et al., 2021a; Cheng et al., 2021; Liu et al., 2021f; Lin et al., 2022b; Shi et al., 2022; Li et al., 2022f; Tang et al., 2022b; Rasoulian et al., 2023; Heidari et al., 2023; Chen & Mo, 2023; Wang et al., 2023f; Chen et al., 2024h) have leveraged Swin Transformer as an encoder, utilizing its features across different vision tasks, as summarized in Table 10.
Table 10: Summary of Methods Applying Knowledge Distillation to the Swin Transformer Model.
@@
MaskFormer (Cheng et al., 2021) Semantic Segmentation
DFR (Zhang et al., 2021a) Weakly Supervised Semantic Segmentation (Scribbles)
SwinNet (Liu et al., 2021f) Salient Object Detection
SwinTrack (Lin et al., 2022b) Siamese Tracking
SwinF (Li et al., 2022f) Target Detection
Swin-UNETR (Tang et al., 2022b) Medical Image Analysis
Ssformer (Shi et al., 2022) Semantic Segmentation
HGI-SAM (Rasoulian et al., 2023) Weakly Supervised Intracranial Hemorrhage Segmentation
HiFormer (Heidari et al., 2023) Medical Image Segmentation
Swin-Fusion (Chen & Mo, 2023) Human Action Recognition
P.Swin (Wang et al., 2023f) Image Classification and Object Detection
SCUNet++ (Chen et al., 2024h) Pulmonary Embolism CT Image Segmentation
@@
SWTformer (Ahmadi & Kasaei, 2024) Weakly Supervised Semantic Segmentation (Class Labels)
6.3 Distillation in Self-supervised learning
@@
Self-Supervised Learning (SSL) is an important field due to its independence from labeled data, and self-distillation is the basis of the most prominent works in SSL, especially in the realm of computer vision. The most prominent work in SSL was based on self-distillation (Chen et al., 2020c). SimCLR (Chen et al., 2020c) proposed a shared network that receives two different augmentations of an image and minimizes their extracted embeddings. Following SimCLR, different extensions have been proposed (He et al., 2020; Chen et al., 2020d; Grill et al., 2020; Misra & Maaten, 2020; Caron et al., 2020; Xu et al., 2021). MOCO (He et al., 2020; Chen et al., 2020d) and BYOL (Grill et al., 2020) are among successful methods where they propose to use the same network but with a different training process. The student is trained using backpropagation, and the teacher is an EMA of the student. SwAP (Caron et al., 2020) uses some prototypes and assigns a code to the augmentations of the same image. The student is then trained to predict the same class ID as the teacher. DINO (Caron et al., 2021) and DINOv2 (Oquab et al., 2023) are the most recent important applications of KD in SSL, where global and local patches of an image pass through similar teacher and student networks. The teacher is updated using EMA, and the representations of the teacher and the student are forced to be similar. This local-to-global approach makes it ideal for segmentation tasks.
@@
Recently, more complex forms of SSL based on KD have been proposed (Misra & Maaten, 2020; Xu et al., 2021; Wang et al., 2023l; Song et al., 2023a; Gao et al., 2022c; Gu et al., 2021a). BINGO (Xu et al., 2021) uses a pre-trained SSL method to group the images of a dataset into a bag of samples. Then, in each epoch, samples are taken from the bag and the inter and intra-sample losses are minimized. Wang et al. (2023l) trains two SSL teachers for images and videos using masked modeling and follows a scenario similar to SimCLR, minimizing the spatial features with the image teacher and the spatio-temporal features with the video teacher. Song et al. (2023a) uses a parallel SSL approach and defines consistency losses between them, and Gao et al. (2022c) integrates a known KD approach into an SSL framework using an SSL pre-trained teacher.
@@
Apart from the mentioned papers, some studies attempt to distill the knowledge of a pre-trained SSL network into a smaller network (Abbasi Koohpayegani et al., 2020; Tejankar et al., 2021; Navaneet et al., 2022; Fang et al., 2021; Dadashzadeh et al., 2022). The gap between two similar networks, one trained fully-supervised and one trained self-supervised, is shown to decrease with a bigger model size (Abbasi Koohpayegani et al., 2020). Navaneet et al. (2022) shows the impact of an MLP layer after representations of the student, while Abbasi Koohpayegani et al. (2020) proposes a similarity-based distillation method for distilling the knowledge of a pre-trained network, which can outperform the corresponding fully-supervised trained network. In a similar approach, SEED (Fang et al., 2021) uses a queue of samples and minimizes the distance of the student’s representation with anchors in the queue and the corresponding distances between the anchors and the teacher’s representation. Dadashzadeh et al. (2022) proposes a novel approach to complement self-supervised pretraining via an auxiliary pretraining phase for small unlabeled datasets.
6.4 Diffusion Distillation
@@
Diffusion models (Ho et al., 2020; Song et al., 2020) generate high-quality images, text, and 3D data through an iterative denoising process that can require hundreds or thousands of steps. This is computationally expensive, limiting deployment on low-resource hardware. KD mitigates this by transferring soft outputs and internal representations from a large teacher model to a smaller student model. The distilled model then achieves the same quality with far fewer sampling steps, requiring significantly less inference time and computation (Luo, 2023).
@@
Several strategies have been developed to accelerate diffusion models. Feature-level methods have the student mimic the teacher’s internal representations (Huang et al., 2023b; Hsiao et al., 2024; Feng et al., 2024b; Zhang et al., 2024b). Sampling process distillation reduces the iterative denoising by either merging multiple steps into one or matching a one-step student’s output to that of the full teacher (Salimans & Ho, 2022; Yin et al., 2024c; Zhou et al., 2024e). Adversarial and data-free approaches align outputs using GAN losses or bootstrapping without requiring large synthetic datasets (Sauer et al., 2024; Mekonnen et al., 2024; Xie et al., 2024b). Consistency models reformulate the denoising into a direct, often single-step transformation, cutting down inference time (Song et al., 2023c; Song & Dhariwal, 2023; Geng et al., 2024; Luo et al., 2023a; Xie et al., 2024a; Lu & Song, 2024). In the following sections, each category is explained in detail.
@@
Feature-level: Feature-level distillation accelerates diffusion model efficiency by bringing the student representations closer to the teacher and reducing inference steps. Recent methods reinforce feature alignment through denoising (Huang et al., 2023b), external guidance (Hsiao et al., 2024), and cross-sample consistency (Feng et al., 2024b), which facilitate convergence and improved sampling efficiency. Furthermore, combining feature distillation with model compression reduces model size and inference time (Zhang et al., 2024b). These techniques accelerate generation without compromising output quality, thereby making diffusion models more realistic.
@@
Sampling Process: Distillation of the sampling procedure is performed to ease the naturally iterative denoising process characteristic of diffusion models. Progressive distillation (Salimans & Ho, 2022) achieves this by consolidating two steps into a single step; through repeated application of this simplification, the duration of the sampling procedure can be drastically shortened. Distribution matching distillation (Yin et al., 2024c) attempts to condense the entire iterative procedure into a single step. This alignment is accomplished by matching the output distribution of a one-step student model directly with that of the combined multi-step teacher model.
@@
Adversarial and Data-free: Adversarial methods (Sauer et al., 2024; Mekonnen et al., 2024) leverage GAN-based losses to enforce the student’s output distribution to closely match that of the teacher. Alternatively, data-free methods like EM distillation (Xie et al., 2024b), employ bootstrapping between successive denoising steps, eliminating the need for large synthetic datasets during training.
@@
Consistency Models: Consistency models (Song et al., 2023c; Song & Dhariwal, 2023; Geng et al., 2024) reformulate the denoising process as a direct, often single transformation from noisy data to its clean version. Their latent-space variants (Luo et al., 2023a; Xie et al., 2024a), along with improvements focusing on stability and scalability (Lu & Song, 2024), achieve significant speedups in inference times. However, training these models may require careful tuning to prevent quality collapse.
@@
Furthermore, recent developments in the area of distillation techniques expanded their accessibility to multi-modal and multi-model settings, thus improving transferability across tasks and diverse architectures. ProlificDreamer (Wang et al., 2023q) and DREAMFUSION (Poole et al., 2022) utilize the methods in the realm of text-to-3D translation, and general-purpose systems like Diff-Instruct (Luo et al., 2023b) can facilitate generalization for the task.
@@
6.5 Visual Recognition Distillation